OpenStudy (anonymous):

let f(x)= 5/x and g(x)=2x^2+5x what two numbers are not in domain of the f o g? A)0,-5/2 B)0,-2/5 c)0,2/5 D)0,5/2 someone help me its a test please!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

can someone please help >.<

OpenStudy (michele_laino):

Please note that you can not divide by zero

OpenStudy (michele_laino):

Please note that your function (fog)(x) is: \[(fog)(x)=f(g(x))=\frac{ 5 }{ g(x) }=\frac{ 5 }{ 2x ^{2}+5x }=\frac{ 5 }{ x(2x+5) }\] note that you can not divide by zero, so?

OpenStudy (michele_laino):

@cx_cassie_

OpenStudy (anonymous):

Thank you